the Laboratory technician needs to make it a 56 L batch of a 40% acid solution. how can the laboratory technician combine a batch of an acid solution that is pure acid with another that is 20% to get the desired concentration?
1
Expert's answer
2019-12-12T11:05:41-0500

1 = 100%

0.2 = 20%


Let x be a volume of 100% solution, y be a volume of 20% solution.

Then

1*x + 0.2*y = 56*0.4,

x+y=56,


1*x + 0.2*y = 22.4,

x= 56-y.


Substituting x=56-y into the first equation

1*(56-y) + 0.2*y = 22.4,


-y +0.2y = -33.6,

-0.8y = -33.6,


0.8y = 33.6,


y=42,


y= 42 L of 20% acid,

x = 56-42,

x= 14 L of 100% acid.

Answer: 14 L of 100% acid, 42 L of 20% acid.